1. Integrate. (a) Vr+3ds (b) x sin (x) cos (2x)dx
Answer #1

a) I- SaS Ju+3 dn Put x+3 =t & (x+3) = d (1) an dt Also a=t-3 Hence, I = ( (2-3) JE dt = Slere- 35e) dt sle - 30) 3S that 3 2

sin (m) cos (2n) dn Sa. (asin(n) cas (2n) du S. (sin (3n) - sinh) an [using 2sin A case Sin( Ar g) +sin (A-) a sinn an 2 Usi